OpenStudy (anonymous):

Ralph runs at a constant rate of 3 miles per hour. How long will it take Ralph to run 15 miles?

OpenStudy (shadowlegendx):

What do you think the answer is?

OpenStudy (shadowlegendx):

@DashL

OpenStudy (gabebae):

He runs 3 miles per hour and he wants to run 15 miles 15 divided by 3 which would equal 5 He would need 5 hours to run 15 miles

OpenStudy (shadowlegendx):

\[\frac{ 3 }{ hour} \rightarrow \frac{ 15 }{ ? }\]

OpenStudy (gabebae):

or you could use the expression 3x=15

OpenStudy (shadowlegendx):

\[\frac{ 3miles }{ hour } \rightarrow \frac{ 15miles }{ ? }\]

OpenStudy (shadowlegendx):

@GabeBae How does that represent the problem

OpenStudy (shadowlegendx):

You should explain how it does, so they can understand it, and apply it to problems in the future.

OpenStudy (gabebae):

Don't worry I'm pretty sure he understands it now

OpenStudy (shadowlegendx):

@DashL Gabe is right, 5 is the answer. But I wish he hadnʻt been so direct about it ;P Do you understand how we got it tho?

OpenStudy (gabebae):

:P sorry

OpenStudy (shadowlegendx):

Itʻs ok, I gave some direct answers when I was new

OpenStudy (shadowlegendx):

But then you learn there is an enjoyment in explaining it out, in hopes of them getting how to do the problem themselves, so they can use those methods in the future. Thus, they have learned

OpenStudy (shadowlegendx):

and our work is done :P

OpenStudy (gabebae):

My work was done 6 minutes ago

OpenStudy (anonymous):

Thanks for the help @GabeBae and @shadowlegendx and the way you get your answer is 3 x what would equal 15 right?

OpenStudy (anonymous):

Oh and @GabeBae I'm a girl ;)

OpenStudy (gabebae):

I'm so sorry I always get slapped by girls irl not even kidding on that one

OpenStudy (anonymous):

@GabeBae It's fine I find it funny every time someone does that :) and Ouch! that must have been painful.

OpenStudy (gabebae):

Nah it felt good :)